You have five(5) 10.0 Farad capacitors. Find the total capacitance if the capacitors are connected in series.

1
Expert's answer
2022-03-21T12:47:41-0400

Assume


"C_1=C_2=C_3=C_4=C_5=10F"

Total capacitance

"\frac{1}{C_t}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}+\frac{1}{C_5}"

"\frac{1}{C_t}=\frac{1}{10}+\frac{1}{10}+\frac{1}{10}+\frac{1}{10}+\frac{1}{10}"

"\frac{1}{C_t}=\frac{5}{10}"

"C_{net}=2F"